The total price including sales tax is $17,551.03. If sales tax rate is 4%, what was price of the van? P=rW P=(4%)17,551.03-702.04=16848.99 wrong,
should be $16875.99

Answer by colliefan(242)

The total price including sales tax is $17,551.03. If sales tax rate is 4%, what was price of the van?

When I think of percents, I always think of the equation
Part = percentage x whole
Work to make sure you realize what the "whole" that the 4% is being applied to is. In this case the percentage would have been multiplied by the cost of the car and then added on to the cost of the car. If x is the cost of the car, the equation you are trying to solve is really:
(x*4%)+x=17551.03
(4%)x+1x=17551.03
(.04)x+1x=17551.03
1.04x=17551.03
x=17551.03/1.04
x=16875.99